OpenStudy (anonymous):

Which of the following is the best estimate for checking the given problem? 478,294,863 x 2,497,832 = 1.19 x 1015

OpenStudy (anonymous):

Use compatible numbers to estimate the following: 4 + 7 = 8 12

OpenStudy (anonymous):

Use compatible numbers to estimate the following: 5,890,361 - 2,043,991 =

OpenStudy (anonymous):

Which of the following is the best estimate for checking the given problem? 3 3 + 3 2 = 8 9